OpenStudy (anonymous):

someone help me out with this question An airport charges a flat rate of $7.95 per day (long term parking). Create a graph of this relationship and evaluate the cost of parking for a car parked for days.

jimthompson5910 (jim_thompson5910):

let x = number of days

OpenStudy (anonymous):

so the answer would be

jimthompson5910 (jim_thompson5910):

if you park for 0 days, ie you don't use the parking service, then you pay nothing so 7.95*0 = 0 is your total cost for x = 0

jimthompson5910 (jim_thompson5910):

if you park for 1 day, then you pay 7.95*1 = 7.95 dollars

jimthompson5910 (jim_thompson5910):

this means x = 1 and y = 7.95

jimthompson5910 (jim_thompson5910):

what happens when x = 2?

OpenStudy (anonymous):

thank you it really helped

jimthompson5910 (jim_thompson5910):

so what is the overall equation?

OpenStudy (anonymous):

the overall equation is 12

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

the overall equation is y = 7.95x to find out how much you pay on a certain day, you plug it into x to find y

jimthompson5910 (jim_thompson5910):

example: day 2, plug in x = 2 and you'll find y

OpenStudy (anonymous):

7.95

jimthompson5910 (jim_thompson5910):

y = 7.95x y = 7.95*2 ... plug in x = 2 y = ??

OpenStudy (anonymous):

15.2????

jimthompson5910 (jim_thompson5910):

7.95*2 = 15.9

jimthompson5910 (jim_thompson5910):

not sure how you got 15.2

jimthompson5910 (jim_thompson5910):

anyways, so you pay 15.90 dollars for 2 days

OpenStudy (anonymous):

ohhh must of type it wrong

OpenStudy (anonymous):

thank you it really helped alot

OpenStudy (anonymous):

what about whats the inverse of the function of f(x) = {(3, 4), (4, 3), (-2, 6)}

jimthompson5910 (jim_thompson5910):

when you have a relation like that, the inverse is simply the set of ordered pairs but each coordinate is swapped

OpenStudy (anonymous):

yes